L^{A}TEX 2ε for class and package writers

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 The L^{A}TEX3 Project

All rights reserved

June 1994

Contents

1 Introduction 1 1.1 Writing classes and packages for L^{A}TEX 2ε . . . . . . . . . . . . . 1 1.2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Further information . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Writing classes and packages 3 2.1 Old versions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Using docstrip and doc . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Is it a class or a package? . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Command names . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Box commands and colour . . . . . . . . . . . . . . . . . . . . . . 4 2.6 Defining text and math characters . . . . . . . . . . . . . . . . . 5 2.7 General style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 The structure of a class or package 8 3.1 Identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.2 Using classes and packages . . . . . . . . . . . . . . . . . . . . . . 9 3.3 Declaring options . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.4 Declarations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.5 Example: a local letter class . . . . . . . . . . . . . . . . . . . . . 12 3.6 Example: a newsletter class . . . . . . . . . . . . . . . . . . . . . 12

4 Commands for class and package writers 13 4.1 Identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2 Loading files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.3 Option declaration . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.4 Within option code commands . . . . . . . . . . . . . . . . . . . 16 4.5 Moving options around . . . . . . . . . . . . . . . . . . . . . . . . 16 4.6 Delaying code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.7 Option processing . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.8 Safe File Commands . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.9 Generating errors . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.10 Defining commands . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.11 Layout parameters . . . . . . . . . . . . . . . . . . . . . . . . . . 22

5 Upgrading L^{A}TEX 2.09 classes and packages 22 5.1 Try it first! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.2 Troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.3 Font commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.4 Obsolete commands . . . . . . . . . . . . . . . . . . . . . . . . . 25

1 Introduction

This document is an introduction to writing classes and packages for L^{A}TEX, with special attention given to upgrading existing L^{A}TEX 2.09 packages to L^{A}TEX 2ε. The latter subject will also be covered in an article by Johannes Braams to be published in TUGboat later this year.

1.1 Writing classes and packages for L^{A}TEX 2ε

L^{A}TEX is a document preparation system that enables the document writer to concentrate on the contents of their text, without bothering too much about the formatting of it. For example, chapters are indicated by \chapter{⟨title⟩} rather than by selecting 18pt bold.

The file that contains the information about how to turn logical structure (like ‘\chapter’) into formatting (like ‘18pt bold ragged right’) is a document class. In addition, some features (such as colour or included graphics) are independent of the document class and these are contained in packages.

One of the largest differences between L^{A}TEX 2.09 and L^{A}TEX 2ε is in the commands used to write packages and classes. In L^{A}TEX 2.09, there was very little support for writing .sty files, and so writers had to resort to using low-level commands.

L^{A}TEX 2ε provides high-level commands for structuring packages. It is also much easier to build classes and packages on top of each other, for example writing a local technical report class based on article.

1.2 Overview

This document contains an overview of how to write classes and packages for L^{A}TEX. It does not introduce all of the commands necessary to write packages: these can be found in either L^{A}TEX: A Document Preparation System or The L^{A}TEX Companion. But it does describe the new commands for structuring classes and packages.

Section 2.7 contains some general advice about writing classes and packages. It describes the difference between classes and packages, the command naming conventions, the use of doc and docstrip, how TEX’s primitive file and box commands interact with L^{A}TEX. It also contains some hints about general L^{A}TEX style.

Section 3 describes the structure of classes and packages. This includes building classes and packages on top of other classes and packages, declaring options and declaring commands. It also contains example classes.

Section 4 lists the new class and package commands.

Section 5 gives detailed advice on how to upgrade existing L^{A}TEX 2.09 classes and packages to L^{A}TEX 2ε.

1.3 Further information

For a general introduction to L^{A}TEX, including the new features of L^{A}TEX 2ε, you should read L^{A}TEX: A Document Preparation System by Leslie Lamport [3].

A more detailed description of the new features of L^{A}TEX, and an overview of more than 150 packages, is to be found in The L^{A}TEX Companion by Michel Goossens, Frank Mittelbach and Alexander Samarin [1].

The L^{A}TEX font selection scheme is based on TEX, which is described in The TEXbook by Donald E. Knuth [2].

There are a number of documentation files which accompany every copy of L^{A}TEX. A copy of L^{A}TEX News will come out with each six-monthly release of L^{A}TEX, and is found in the files ltnews*.tex. The author’s guide L^{A}TEX 2ε for Authors describes the new L^{A}TEX document features; it is in usrguide.tex. The guide L^{A}TEX 2ε Font Selection describes the L^{A}TEX font selection scheme for class- and package-writers; it is in fntguide.tex.

We are gradually turning the source code for L^{A}TEX into a L^{A}TEX document L^{A}TEX: the program. This document includes an index of L^{A}TEX commands and can be typeset from source2e.tex.

For more information about TEX and L^{A}TEX, please contact your local TEX Users Group, or the international TEX Users Group, P. O. Box 869, Santa Barbara, CA 93102-0869, USA, Fax: +1 805 963 8358, E-mail: tug@tug.org.

2 Writing classes and packages

This section covers some general points concerned with writing L^{A}TEX classes and packages.

2.1 Old versions

If you are upgrading an existing L^{A}TEX 2.09 style file then we recommend freezing the 2.09 version and no longer maintaining it. Experience with the various dialects of L^{A}TEX which existed in the early 1990’s suggests that maintaining packages for different versions of L^{A}TEX is almost impossible. It will, of course, be necessary for some organisations to maintain both versions in parallel for some time but this is not essential for those packages and classes supported by individuals: they should support only the new standard L^{A}TEX 2ε, not obsolete versions of L^{A}TEX.

2.2 Using docstrip and doc

If you are going to write a large class or package for L^{A}TEX then you should consider using the doc software which comes with L^{A}TEX. L^{A}TEX classes and packages written using this can be processed in two ways: they can be run through L^{A}TEX, to produce documentation; and they can be processed with docstrip, to produce the .cls or .sty file.

The doc software can automatically generate indexes of definitions, indexes of command use, and change-log lists. It is very useful for maintaining and documenting large TEX sources.

The documented sources of the L^{A}TEX kernel itself ,and of the standard classes, etc, are doc documents; they are in the .dtx files in the distribution. You can, in fact, typeset the source code of the kernel as one long document, complete with index, by running L^{A}TEX on source2e.tex.

For more information on doc and docstrip, consult the files docstrip.dtx, doc.dtx, and The L^{A}TEX Companion. For examples of its use, look at the .dtx files.

2.3 Is it a class or a package?

The first thing to do when you want to put some new L^{A}TEX commands in a file is to decide whether it should be a document class or a package. The rule of thumb is:

if the commands could be used with any document class, then make them a package; and if not, then make them a class.

There are two major types of class: those like article, report or letter, which are free-standing; and those which are extensions or variations of other classes—for example, the proc document class, which is built on the article document class.

Thus, a company might have a local ownlet class for printing letters with their own headed notepaper. Such a class would build on top of the existing letter class but it cannot be used with any other document class, so we

have ownlet.cls rather than ownlet.sty.

The graphics package, in contrast, provides commands for including images into a L^{A}TEX document. Since these commands can be used with any document

class, we have graphics.sty rather than graphics.cls.

2.4 Command names

L^{A}TEX has three types of command.

There are the author commands, such as \section, \emph and \times: most of these have short names, all in lower case.

There are the class and package writer commands: most of these have long mixed-case names such as the following.

   \InputIfFileExists  \RequirePackage  \PassOptionsToClass

Finally, there are the internal commands used in the L^{A}TEX implementation, such as \@tempcnta, \@ifnextchar and \@eha: most of these commands contain @ in their name, which means they cannot be used in documents, only in class and package files.

Unfortunately, for historical reasons the distinction between these commands is often blurred. For example, \hbox is an internal command which should only be used in the L^{A}TEX kernel, whereas \m@ne is the constant −1 and could have

been \MinusOne.

However, this rule of thumb is still useful: if a command has @ in its name then it is not part of the supported L^{A}TEX language—and its behaviour may change in future releases! If a command is mixed-case, or is described in L^{A}TEX: A Document Preparation System, then you can rely on future releases of L^{A}TEX 2ε supporting the command.

2.5 Box commands and colour

Even if you do not intend to use colour in your own documents, by taking note of the points in this section you can ensure that your class or package is compatible with the color package. This may benefit people using your class or package who have access to colour printers.

The simplest way to ensure ‘colour safety’ is to always use L^{A}TEX box commands rather than TEX primitives, that is use \sbox rather than \setbox, \mbox rather than \hbox and \parbox or \minipage rather than \vbox. The L^{A}TEX box commands have new options which mean that they are now as powerful as the TEX primitives.

As an example of what can go wrong, consider that in {\ttfamily ⟨text⟩} the font is restored just before the }, whereas in the similar looking construction {\color{green} ⟨text⟩} the colour is restored just after the final }. Normally this distinction does not matter at all; but consider a primitive TEX box assignment such as:

   \setbox0=\hbox{\color{green} ⟨text⟩}

Now the colour-restore occurs after the } and so is not stored in the box. Exactly what bad effects this can have depends on how colour is implemented: it can range from getting the wrong colours in the rest of the document, to causing errors in the dvi-driver used to print the document.

Also of interest is the command \normalcolor. This is normally just \relax (i.e., does nothing) but you can use it rather like \normalfont to set regions of the page such as captions or section headings to the ‘main document colour’.

2.6 Defining text and math characters

Because L^{A}TEX 2ε supports different encodings, definitions of commands for producing symbols, accents, composite glyphs, etc. must be defined using the commands provided for this purpose and described in L^{A}TEX 2ε Font Selection. This part of the system is still under development so such tasks should be undertaken with great caution.

Also, \DeclareRobustCommand should be used for encoding-independent commands of this type.

Note that it is no longer possible to refer to the math font set-up outside math mode: for example, neither \textfont 1 nor \scriptfont 2 are guaranteed to be defined in other modes.

2.7 General style

The new system provides many commands designed to help you produce wellstructured class and package files that are both robust and portable. This section outlines some ways to make intelligent use of these.

2.7.1 Loading other files

L^{A}TEX provides the commands \LoadClass and \RequirePackage for using classes or packages inside other classes or packages. We recommend strongly that you use them, rather than the primitive \input command, for a number of reasons.

Files loaded with \input ⟨filename⟩ will not be listed in the \listfiles list.

If a package is always loaded with \RequirePackage or \usepackage then, even if its loading is requested several times, it will be loaded only once. By contrast, if it is loaded with \input then it can be loaded more than once; such an extra loading may waste time and memory and it may produce strange results.

If a package provides option-processing then, again, strange results are possible if the package is \input rather than loaded by means of \usepackage or

\RequirePackage.

If the package foo.sty loads the package baz.sty by use of \input baz.sty then the user will get a warning:

   LaTeX Warning: You have requested package `foo',
                  but the package provides `baz'.

Thus, for several reasons, using \input to load packages is not a good idea.

Unfortunately, if you are upgrading the file myclass.sty to a class file then you have to make sure that any old files which contain \input myclass.sty still work.

This is also true of the standard classes (article, book and report), since a lot of existing L^{A}TEX 2.09 document styles contain \input article.sty. The approach which we took was to provide minimal files article.sty, book.sty and report.sty, which load the appropriate class files.

For example, article.sty contains just the following lines:

   \NeedsTeXFormat{LaTeX2e}
   \@obsoletefile{article.sty}{article.cls}
   \LoadClass{article}

You may wish to do the same or, if you think that it is safe to do so, you may decide to just remove myclass.sty.

2.7.2 Make it robust

We consider it good practice, when writing packages and classes, to use L^{A}TEX commands as much as possible.

Thus, instead of using \def... we recommend using one of \newcommand,

\renewcommand or \providecommand; \CheckCommand is also useful. Doing this

makes it less likely that you will inadvertently redefine a command, giving unexpected results.

When you define an environment, use \newenvironment or \renewenvironment
instead \def\foo{...} and \def\endfoo{...}.

If you need to set or change the value of a ⟨dimen⟩ or ⟨skip⟩ register, use

\setlength.

To manipulate boxes, use L^{A}TEX commands such as \sbox, \mbox and \parbox

rather than \setbox, \hbox and \vbox.
Use \PackageError, \PackageWarning or \PackageInfo (or the equivalent

class commands) rather than \@latexerr, \@warning or \wlog.

It is still possible to declare options by defining \ds@⟨option⟩ and calling

\@options; but we recommend the \DeclareOption and \ProcessOptions

commands instead. These are more powerful and use less memory. So rather than using:

   \def\ds@draft{\overfullrule 5pt}
   \@options

you should use:

   \DeclareOption{draft}{\setlength{\overfullrule}{5pt}}
   \ProcessOptions

The advantage of this kind of practice is that your code is more readable and, more important, that it is less likely to break when used with future versions of L^{A}TEX.

2.7.3 Make it portable

It is also sensible to make your files are as portable as possible. To ensure this; they should contain only visible 7-bit text; and the filenames should contain at most eight characters (plus the three letter extension).

It is also useful if local classes or packages have a common prefix, for example the University of Nowhere classes might begin with unw. This helps to avoid every University having its own thesis class, all called thesis.cls.

If you rely on some features of the L^{A}TEX kernel, or on a package, please specify the release-date you need. For example, the package error commands were introduced in the June 1994 release so, if you use them then you should put:

   \NeedsTeXFormat{LaTeX2e}[1994/06/01]

2.7.4 Useful hooks

Some packages and document styles had to redefine the command \document or \enddocument to achieve their goal. This is no longer necessary. You can now use the ‘hooks’ \AtBeginDocument and \AtEndDocument (see Section 4.6). Again, using these hooks makes it less likely that your code breaks with future versions of L^{A}TEX. It also makes it more likely that your package can work together with someone else’s.

3 The structure of a class or package

L^{A}TEX 2ε classes and packages have more structure than L^{A}TEX 2.09 style files did. The outline of a class or package file is:

Identification The file says that it is a L^{A}TEX 2ε package or class, and gives a short description of itself.

Preliminary declarations Here the file declares some commands and can also load other files. Usually these commands will be just those needed for the code used in the declared options.

Options The file declares and processes its options.

More declarations This is where the file does most of its work: declaring new variables, commands and fonts; and loading other files.

3.1 Identification

The first thing a class or package file does is identify itself. Package files do this as follows:

   \NeedsTeXFormat{LaTeX2e}

\ProvidesPackage{⟨package⟩}[⟨date⟩ ⟨other information⟩]

For example:

   \NeedsTeXFormat{LaTeX2e}
   \ProvidesPackage{latexsym}[1994/06/01 Standard LaTeX package]

Class files do this as follows:

   \NeedsTeXFormat{LaTeX2e}

\ProvidesClass{⟨class-name⟩}[⟨date⟩ ⟨other information⟩]

For example:

   \NeedsTeXFormat{LaTeX2e}
   \ProvidesClass{article}[1994/06/01 Standard LaTeX class]

The ⟨date⟩ should be given in the form ‘yyyy/mm/dd’. This date is checked whenever a user specifies a date in their \documentclass or \usepackage command. For example, if you wrote:

   \documentclass{article}[1995/12/23]

then users at a different location would get a warning that their copy of article was out of date.

The description of a class is displayed when the class is used. The description of a package is put into the log file. These descriptions are also displayed by the

\listfiles command.

3.2 Using classes and packages

The first major difference between L^{A}TEX 2.09 style files and L^{A}TEX 2ε packages and classes is that L^{A}TEX 2ε supports modularity, in the sense of building files from small building-blocks rather than using large single files.

A L^{A}TEX package or class can load a package as follows:

\RequirePackage[⟨options⟩]{⟨package⟩}[⟨date⟩]

For example:

   \RequirePackage{ifthen}[1994/06/01]

This command has the same syntax as the author command \usepackage. It allows packages or classes to use features provided by other packages. For example, by loading the ifthen package, a package writer can use the ‘if… then… else… ’ commands provided by that package.

A L^{A}TEX class can load one other class as follows:

\LoadClass[⟨options⟩]{⟨class-name⟩}[⟨date⟩]

For example:

   \LoadClass[twocolumn]{article}

This command has the same syntax as the author command \documentclass. It allows classes to be based on the syntax and appearance of another class. For example, by loading the article class, a class writer only has to change the bits of article they don’t like, rather than writing a new class from scratch.

3.3 Declaring options

The other major difference between L^{A}TEX 2.09 styles and L^{A}TEX 2ε packages and classes is in option handling. Packages and classes can now declare options and these can be specified by authors; for example, the twocolumn option is declared by the article class.

An option is declared as follows:

   \DeclareOption{⟨option⟩}{⟨code⟩}

For example, the dvips option to the graphics package is implemented as:

   \DeclareOption{dvips}{\input{dvips.def}}

This means that when an author writes \usepackage[dvips]{graphics}, the file dvips.def is loaded. As another example, the a4paper option is declared

in the article class to set the \paperheight and \paperwidth lengths:
   \DeclareOption{a4paper}{%
      \setlength{\paperheight}{297mm}%
      \setlength{\paperwidth}{210mm}%
   }

Sometimes a user will request an option which the class or package has not explicitly declared. By default this will produce a warning (for classes) or error (for packages); this behaviour can be altered as follows:

   \DeclareOption*{⟨code⟩}

For example, to make the package fred produce a warning rather than an error for unknown options, you could specify:

   \DeclareOption*{%
      \PackageWarning{fred}{Unknown option `\CurrentOption'}%
   }

Then, if an author writes \usepackage[foo]{fred}, they will get a warning

Package fred Warning:  Unknown option `foo'. As another example, the

fontenc package tries to load a file ⟨ENC ⟩enc.def whenever the ⟨ENC ⟩ option is used. This can be done by writing:

   \DeclareOption*{%
      \input{\CurrentOption enc.def}%
   }

It is possible to pass options on to another package or class, using the command

\PassOptionsToPackage or \PassOptionsToClass. For example, to pass every

unknown option on to the article class, you can use:

   \DeclareOption*{%
      \PassOptionsToClass{\CurrentOption}{article}%
   }

If you do this then you should make sure you load the class at some later point, otherwise the options will never be processed!

So far, we have explained only how to declare options, not how to execute them. To process the options with which the file was called, you should use:

   \ProcessOptions

This executes the ⟨code⟩ for each option that was both specified and declared (see Section 4.7 for details of how this is done).

For example, if the jane package file contains:

   \DeclareOption{foo}{\typeout{Saw foo.}}
   \DeclareOption{baz}{\typeout{Saw baz.}}
   \DeclareOption*{\typeout{What's \CurrentOption?}}
   \ProcessOptions

and an author writes \usepackage[foo,bar]{jane}, then they will see the

messages Saw foo. and What's bar?

3.4 Declarations

Most of the work of a class or package is in defining new commands, or changing the appearance of documents. This is done in the body of the package, using

commands such as \newcommand or \setlength.

L^{A}TEX 2εprovides several new commands to help class and package writers; these are described in detail in Section 4.

There are three things that every class file must contain: these are a definition

of \normalsize and values for \textwidth and \textheight. So a minimal

document class file is:

   \NeedsTeXFormat{LaTeX2e}
   \ProvidesClass{minimal}[1994/04/01 Minimal class]
   \renewcommand{\normalsize}{\fontsize{10}{12}\selectfont}
   \setlength{\textwidth}{6.5in}
   \setlength{\textheight}{8in}

However, this class file will not support footnotes, marginals, floats, etc. nor will it provide any of the 2-letter font commands such as \rm; thus most classes will contain more than this minimum!

3.5 Example: a local letter class

A company may have its own letter class, for setting letters in the company style. This section shows a simple implementation of such a class, although a real class would need more structure.

The class begins by announcing itself as neplet.cls.

   \NeedsTeXFormat{LaTeX2e}
   \ProvidesClass{neplet}[1995/04/01 NonExistent Press letter class]

Then this next bit passes any options on to the letter class, which is loaded with the a4paper option.

   \DeclareOption*{\PassOptionsToClass{\CurrentOption}{letter}}
   \ProcessOptions
   \LoadClass[a4paper]{letter}

In order to use the company letter head, it redefines the firstpage page style: this is the page style that is used on the first page of letters.

   \renewcommand{\ps@firstpage}{%
      \renewcommand{\@oddhead}{⟨letterhead goes here⟩}%
      \renewcommand{\@oddfoot}{⟨letterfoot goes here⟩}%
   }

And that’s it!

3.6 Example: a newsletter class

A simple newsletter can be typeset with L^{A}TEX, using a variant of the article class. The class begins by announcing itself as smplnews.cls.

   \NeedsTeXFormat{LaTeX2e}
   \ProvidesClass{smplnews}[1995/04/01 The Simple News newsletter class]
   \newcommand{\headlinecolor}{\normalcolor}

It passes most specified options on to the article class: apart from the onecolumn option, which is switched off, and the green option, which sets the headline in green.

   \DeclareOption{onecolumn}{\OptionNotUsed}
   \DeclareOption{green}{\renewcommand{\headlinecolor}{\color{green}}}
   \DeclareOption*{\PassOptionsToClass{\CurrentOption}{article}}
   \ProcessOptions

It then loads the class article with the option twocolumn.

   \LoadClass[twocolumn]{article}

Since the newsletter is to be printed in colour, it now loads the color package. The class does not specify a device driver option since this should be specified by the user of the smplnews class.

   \RequirePackage{color}

The class then redefines \maketitle to produce the title in 72pt Helvetica bold oblique, in the appropriate colour.

   \renewcommand{\maketitle}{%
      \twocolumn[%
         \fontsize{72}{80}\fontfamily{phv}\fontseries{b}%
         \fontshape{sl}\selectfont\headlinecolor
         \@title
      ]%
   }

It redefines \section and switches off section numbering.

   \renewcommand{\section}{%
      \@startsection
         {section}{1}{0pt}{-1.5ex plus -1ex minus -.2ex}%
         {1ex plus .2ex}{\large\sffamily\slshape\headlinecolor}%
   }
   \setcounter{secnumdepth}{0}

It also sets the three essential things.

   \renewcommand{\normalsize}{\fontsize{9}{10}\selectfont}
   \setlength{\textwidth}{17.5cm}
   \setlength{\textheight}{25cm}

In practice, a class would need more than this: it would provide commands for issue numbers, authors of articles, page styles and so on; but this skeleton gives a start. The ltnews class file is not much more complex than this one.

4 Commands for class and package writers

This section describes briefly each of the new commands for class and package writers. To find out about other aspects of the new system, you should also read L^{A}TEX: A Document Preparation System, The L^{A}TEX Companion and L^{A}TEX 2ε for Authors.

4.1 Identification

The first group of commands discussed here are those used to identify your class or package file.

\NeedsTeXFormat {⟨format-name⟩} [⟨release-date⟩]

This command tells TEX that this file should be processed using a format with name ⟨format-name⟩. You can use the optional argument ⟨release-date⟩ to further specify the earliest release date of the format that is needed. When the release date of the format is older than the one specified a warning will be generated. The standard ⟨format-name⟩ is LaTeX2e. The date, if present, must be in the form yyyy/mm/dd. Example:

   \NeedsTeXFormat{LaTeX2e}[1994/06/01]

\ProvidesClass {⟨class-name⟩} [⟨release-info⟩] \ProvidesPackage {⟨package-name⟩} [⟨release-info⟩]

This declares that the current file contains the definitions for the document class ⟨class-name⟩ or package ⟨package-name⟩.

The optional ⟨release-info⟩, if used, must contain:

• the release date of this version of the file, in the form yyyy/mm/dd;

• optionally followed by a space and a short description, possibly including a version number.

The above syntax should be followed exactly so that this information can be

used by \LoadClass or \documentclass (for classes) or \RequirePackage or

\usepackage (for packages) to test that the release is not too old.

The whole of this ⟨release-info⟩ information is displayed by \listfiles and should therefore not be too long. Example:

   \ProvidesClass{article}[1994/06/01 v1.0 Standard LaTeX class]
   \ProvidesPackage{ifthen}[1994/06/01 v1.0 Standard LaTeX package]

\ProvidesFile {⟨file-name⟩} [⟨release-info⟩]

This is similar to the two previous commands except that here the full filename, including the extension, must be given. It is used for declaring any files other than main class and package files. Example:

   \ProvidesFile{T1enc.def}[1994/06/01 v1.0 Standard LaTeX file]

4.2 Loading files

This group of commands can be used to create your own document class or package by building on existing classes or packages.

\RequirePackage [⟨options-list⟩] {⟨package-name⟩} [⟨release-info⟩]

Packages and classes should use this command to load other packages. Its use is the same as the author command \usepackage.

Example:

   \RequirePackage{ifthen}[1994/06/01]

\LoadClass [⟨options-list⟩] {⟨class-name⟩} [⟨release-info⟩]

This command is for use only in class files, it cannot be used in packages files. It is used in the same way as \documentclass, to load a class file.

Also, it can be used at most once within a class file.

Example:

   \LoadClass{article}[1994/06/01]

4.3 Option declaration

The following commands deal with the declaration and handling of options to document classes and packages.

There are some commands designed especially for use within the ⟨code⟩ argument of these commands (see below).

\DeclareOption {⟨option-name⟩} {⟨code⟩}

This makes ⟨option-name⟩ a ‘declared option’ of the class or package in which it is put.

The ⟨code⟩ argument contains the code to be executed if that option is specified for the class or package; it can contain any valid L^{A}TEX 2ε construct.

Example:

   \DeclareOption{twoside}{\@twosidetrue}

\DeclareOption* {⟨code⟩}

This declares the ⟨code⟩ to be executed for every option which is specified for, but otherwise not explicitly declared by, the class or package; this code is called the ‘default option code’ and it can contain any valid L^{A}TEX 2ε construct.

If a class file contains no \DeclareOption* then, by default, all specified but undeclared options for that class will be silently passed to all packages (as will the specified and declared options for that class).

If a package file contains no \DeclareOption* then, by default, each specified but undeclared option for that package will produce an error.

4.4 Within option code commands

These two commands can be used only within the ⟨code⟩ argument of either \DeclareOption or \DeclareOption*. Other commands commonly used within these arguments can be found in the next few subsections.

\CurrentOption

This expands to the name of the current option.

\OptionNotUsed

This causes the current option to be added to the list of ‘unused options’.

4.5 Moving options around

These two commands are also very useful within the ⟨code⟩ argument of

\DeclareOption or \DeclareOption*.

\PassOptionsToPackage {⟨options-list⟩} {⟨package-name⟩} \PassOptionsToClass {⟨options-list⟩} {⟨class-name⟩}

The command \PassOptionsToPackage passes the options in ⟨options-list⟩ to package ⟨package-name⟩. This means that it adds the ⟨option-list⟩ to the list of options used by any future \RequirePackage or \usepackage command for package ⟨package-name⟩.

Example:

   \PassOptionsToPackage{foo,bar}{fred}
   \RequirePackage[baz]{fred}

is the same as:

   \RequirePackage[foo,bar,baz]{fred}

Similarly, \PassOptionsToClass may be used in a class file to pass options to another class to be loaded with \LoadClass.

4.6 Delaying code

These first two commands are also intended primarily for use within the ⟨code⟩

argument of \DeclareOption or \DeclareOption*.
\AtEndOfClass {⟨code⟩}
\AtEndOfPackage {⟨code⟩}

These commands declare ⟨code⟩ that is saved away internally and then executed after processing the whole of the current class or package file.

Repeated use of these commands is permitted: the code in the arguments is stored (and later executed) in the order of their declarations.

\AtBeginDocument {⟨code⟩}
\AtEndDocument {⟨code⟩}

These commands declare ⟨code⟩ to be saved internally and executed while L^{A}TEX

is executing \begin{document} or \end{document}.

The ⟨code⟩ specified in the argument to \AtBeginDocument is executed near the end of the \begin{document} code, after the font selection tables have been set up. It is therefore a useful place to put code which needs to be executed after everything has been prepared for typesetting and when the normal font for the document is the current font.

The ⟨code⟩ specified in the argument to \AtEndDocument is executed at the beginning of the \end{document} code, before the final page is finished and before any leftover floating environments are processed. If some of the ⟨code⟩ is to be executed after these two processes, you should include a \clearpage at the appropriate point in ⟨code⟩.

Repeated use of these commands is permitted: the code in the arguments is stored (and later executed) in the order of their declarations.

4.7 Option processing

This command executes the ⟨code⟩ for each selected option.

\ProcessOptions

We shall first describe how \ProcessOptions works in a package file, and then how this differs in a class file.

To understand in detail what \ProcessOptions does in a package file, you have to know the difference between local and global options.

• Local options are those which have been explicitly specified for this particular package in the ⟨options⟩ argument of any of these:

         \PassOptionsToPackage{⟨options⟩} \usepackage[⟨options⟩]
         \RequirePackage[⟨options⟩]

• Global options are any other options that are specified by the author in the ⟨options⟩ argument of \documentclass[⟨options⟩].

For example, suppose that a document begins:

   \documentclass[german,twocolumn]{article}
   \usepackage{gerhardt}

whilst package gerhardt calls package fred with:

   \PassOptionsToPackage{german,dvips,a4paper}{fred}
   \RequirePackage[errorshow]{fred}

then:

   • fred’s local options are german, dvips, a4paper and errorshow;

• fred’s only global option is twocolumn.

When \ProcessOptions is called, the following happen.

• First, for each option so far declared in fred.sty by \DeclareOption, it looks to see if that option is either a global or a local option for fred: if it is then the corresponding code is executed.

This is done in the order in which these options were declared in fred.sty.

• Then, for each remaining local option, the command \ds@⟨option⟩ is executed if it has been defined somewhere (other than by a \DeclareOption); otherwise, the ‘default option code’ is executed. If no default option code has been declared then an error message is produced.

This is done in the order in which these options were specified.

Throughout this process, the system ensures that the code declared for an option is executed at most once.

Returning to the example, if fred.sty contains:

   \DeclareOption{dvips}{\typeout{DVIPS}}
   \DeclareOption{german}{\typeout{GERMAN}}
   \DeclareOption{french}{\typeout{FRENCH}}
   \DeclareOption*{\PackageWarning{fred}{Unknown `\CurrentOption'}}
   \ProcessOptions

then the result of processing this document will be:

   DVIPS
   GERMAN
   Package fred Warning: Unknown `a4paper'.
   Package fred Warning: Unknown `errorshow'.

Note the following:

• the code for the dvips option is executed before that for the german option, because that is the order in which they are declared in fred.sty;

• the code for the german option is executed only once, when the declared options are being processed;

• the a4paper and errorshow options produce the warning from the code declared by \DeclareOption* (in the order in which they were specified), whilst the twocolumn option does not: this is because twocolumn is a global option.

In a class file, \ProcessOptions works in the same way, except that: all options are local; and the default value for \DeclareOption* is \OptionNotUsed rather than an error.

\ProcessOptions* \@options

This is like \ProcessOptions but it executes the options in the order specified in the calling commands, rather than in the order of declaration in the class or package. For a package this means that the global options are processed first.

The \@options command from L^{A}TEX 2.09 has been made equivalent to this in order to ease the task of updating old document styles to L^{A}TEX 2ε class files.

\ExecuteOptions {⟨options-list⟩}

For each option in the ⟨options-list⟩, in order, this command simply executes the command \ds@⟨option⟩ (if this command is not defined, then that option is silently ignored). It can be used to provide a ‘default option list’ just before \ProcessOptions. For example, suppose that in a class file you want to set up the default design to be: two-sided printing; 11pt fonts; in two columns. Then it could specify:

  \ExecuteOptions{11pt,twoside,twocolumn}

4.8 Safe File Commands

These commands deal with file input; they ensure that the non-existence of a requested file can be handled in a user-friendly way.

\IfFileExists {⟨file-name⟩} {⟨true⟩} {⟨false⟩}

If the file exists then the code specified in ⟨true⟩ is executed.

If the file does not exist then the code specified in ⟨false⟩ is executed.

This command does not input the file.

\InputIfFileExists {⟨file-name⟩} {⟨true⟩} {⟨false⟩}

This inputs the file ⟨file-name⟩ if it exists and, immediately before the input, the code specified in ⟨true⟩ is executed.

If the file does not exist then the code specified in ⟨false⟩ is executed.

It is implemented using \IfFileExists.

4.9 Generating errors

These commands should be used by third party classes and packages to report errors, or to provide information to authors.

\ClassError {⟨class-name⟩} {⟨error-text⟩} {⟨help-text⟩} \PackageError {⟨package-name⟩} {⟨error-text⟩} {⟨help-text⟩}

These produce an error message. The ⟨error-text⟩ is displayed and the ? error prompt is shown. If the user types h, they will be shown the ⟨help-text⟩.

Within the ⟨error-text⟩ and ⟨help-text⟩: \protect can be used to stop a command from expanding; \MessageBreak causes a line-break; and \space prints a space.

Note that the ⟨error-text⟩ will have a full stop added to it, so do not put one into the argument.

For example:

   \newcommand{\foo}{FOO}
   \PackageError{ethel}{%
      Your hovercraft is full of eels,\MessageBreak
      and \protect\foo\space is \foo
   }{%
      Oh dear! Something's gone wrong.\MessageBreak
      \space \space Try typing \space <return>

      \space to proceed, ignoring \protect\foo.
   }

produces this display:

   ! Package ethel Error: Your hovercraft is full of eels,
   (ethel)                and \foo is FOO.
   See the ethel package documentation for explanation.

If the user types h, this will be shown:

   Oh dear! Something's gone wrong.
     Try typing  <return>  to proceed, ignoring \foo.

\ClassWarning {⟨class-name⟩} {⟨warning-text⟩} \PackageWarning {⟨package-name⟩} {⟨warning-text⟩} \ClassWarningNoLine {⟨class-name⟩} {⟨warning-text⟩} \PackageWarningNoLine {⟨package-name⟩} {⟨warning-text⟩} \ClassInfo {⟨class-name⟩} {⟨info-text⟩} \PackageInfo {⟨package-name⟩} {⟨info-text⟩}

The four Warning commands are similar to the error commands, except that they produce only a warning on the screen, with no error prompt.

The first two, Warning versions, also show the line number where the warning occurred, whilst the second two, WarningNoLine versions, do not.

The two Info commands are similar except that they log the information only in the transcript file, including the line number. There are no NoLine versions of these two.

Within the ⟨warning-text⟩ and ⟨info-text⟩: \protect can be used to stop a command from expanding; \MessageBreak causes a line-break; and \space prints a space. Also, these should not end with a full stop as one is automatically added.

4.10 Defining commands

L^{A}TEX 2ε provides some extra methods of (re)defining commands that are intended for use in class and package files.

\DeclareRobustCommand {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩}

This command takes the same arguments as \newcommand but it declares a robust command, even if some code within the⟨definition⟩ is fragile. You can use this command to define new robust commands, or to redefine existing commands and make them robust. A log is put into the transcript file if a command is redefined.

For example, if \seq is defined as follows:

   \DeclareRobustCommand{\seq}[2][n]{%
     \ifmmode
       #1_{1}\ldots#1_{#2}%
     \else
       \PackageWarning{fred}{You can't use \protect\seq\space in text}%
     \fi
   }

Then the command \seq can be used in moving arguments, even though \ifmmode cannot, for example:

   \section{Stuff about sequences $\seq{x}$}

Note also that there is no need to put a \relax before the \ifmmode at the beginning of the definition; this is because the protection given by this \relax against expansion at the wrong time will be provided internally.

\CheckCommand {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩}

This takes the same arguments as \newcommand but, rather than define ⟨cmd⟩, it just checks that the current definition of ⟨cmd⟩ is exactly as given by ⟨definition⟩. An error is raised if these definitions differ.

This command is useful for checking the state of the system before your package starts altering the definitions of commands. It allows you to check, in particular, that no other package has redefined the same command.

4.11 Layout parameters

\paperheight
\paperwidth

These two parameters are usually set by the class to be the size of the paper being used. This should be actual paper size, unlike \textwidth and \textheight which are the size of the main text body within the margins.

5 Upgrading L^{A}TEX 2.09 classes and packages

This section describes the changes you may need to make when you upgrade an existing L^{A}TEX style to a package or class but we shall start in optimistic mode.

Many existing style files will run with L^{A}TEX 2ε without any modification to the file itself. When everything is running OK, please put a note in the newly created package or class file to record that it runs with the new standard L^{A}TEX; then distribute it to your users.

5.1 Try it first!

The first thing you should do is to test your style in ‘compatibilty mode’. The only change you need to make in order to do this is, possibly, to change the extension of the file to .cls: you should make this change only if your file was used as a main document style. Now, without any other modifications, run L^{A}TEX 2ε on a document that uses your file. This assumes that you have a suitable collection of files that tests all the functionality provided by your style file. (If you haven’t, now is the time to make one!)

You now need to change the test document files so that they are L^{A}TEX 2ε documents: see L^{A}TEX 2ε for Authors for details of how to do this and then try them again. You have now tried the test documents in both L^{A}TEX 2ε native mode and L^{A}TEX 2.09 compatibility mode.

5.2 Troubleshooting

If your file does not work with L^{A}TEX 2ε, there are two likely reasons.

• L^{A}TEX now has a robust, well-defined designer’s interface for selecting fonts, which is very different from the L^{A}TEX 2.09 internals.

• Your style file may have used some L^{A}TEX 2.09 internal commands which have changed, or which have been removed.

When you are debugging your file, you will probably need more information than is normally displayed by L^{A}TEX 2ε. This is achieved by resetting the counter errorcontextlines from its default value of −1 to a much higher value, e.g. 999.

5.3 Font commands

Some font and size commands are now defined by the document class rather than by the L^{A}TEX kernel. If you are upgrading a L^{A}TEX 2.09 document style to a class that does not load one of the standard classes, then you will probably need to add definitions for these commands.

\rm \sf \tt \bf \it \sl \sc

None of these short-form font selection commands are defined in the L^{A}TEX 2ε kernel. They are defined by all the standard class files.

If you want to define them in your class file, there are several reasonable ways to do this.

One possible definition is:

   \newcommand{\rm}{\rmfamily}
   ...
   \newcommand{\sc}{\scshape}

This would make the font commands orthogonal; for example {\bf\it text} would produce bold italic, thus: text. It will also make them produce an error if used in math mode.

Another possible definition is:

   \DeclareOldFontCommand{\rm}{\rmfamily}{\mathrm}
   ...
   \DeclareOldFontCommand{\sc}{\scshape}{\mathsc}

This will make \rm act like \rmfamily in text mode (see above) and it will make \rm select the \mathrm math alphabet in math mode.

Thus ${\rm math} = X + 1$ will produce ‘math = X + 1’.

If you do not want font selection to be orthogonal then you can follow the standard classes and define:

   \DeclareOldFontCommand{\rm}{\normalfont\rmfamily}{\mathrm}
   ...
   \DeclareOldFontCommand{\sc}{\normalfont\scshape}{\mathsc}

This means, for example, that {\bf\it text} will produce medium weight (rather than bold) italic, thus: text.

\normalsize
\@normalsize

The command \@normalsize is retained for compatibility with L^{A}TEX 2.09 packages which may have used its value; but redefining it in a class file will have no effect since it is always reset to have the same meaning as \normalsize.

This means that classes must now redefine \normalsize rather than redefining \@normalsize; for example (a rather incomplete one):

   \renewcommand{\normalsize}{\fontsize{10}{12}\selectfont}

Note that \normalsize is defined by the L^{A}TEX kernel to be an error message.

\tiny \footnotesize \small \large
\Large \LARGE \huge \Huge

None of these other ‘standard’ size-changing commands are defined in the kernel: each needs to be defined in a class file if you need it. They are all defined by the standard classes.

This means you should use \renewcommand for \normalsize and \newcommand

for the other size-changing commands.

5.4 Obsolete commands

Some packages will not work with L^{A}TEX 2ε, normally because they relied on an internal L^{A}TEX command which was never supported and has now changed, or been removed.

In many cases there will now be a robust, high-level means of achieving what previously required low-level commands. Please consult Section 4 to see if you can now use the L^{A}TEX 2ε class and package writers commands.

Also, of course, if your package or class redefined any of the kernel commands (i.e. those defined in the files latex.tex, slitex.tex, lfonts.tex, sfonts.tex) then you will need to check it very carefully against the new kernel in case the implementation has changed or the command no longer exists in the L^{A}TEX2e kernel.

Too many of the internal commands of L^{A}TEX 2.09 have been re-implemented or removed to be able to list them all here. You must check any that you have used or changed.

We shall, however, list some of the more important commands which are no longer supported.

\tenrm \elvrm \twlrm . . .
\tenbf \elvbf \twlbf . . .
\tensf \elvsf \twlsf . . .

⋮

The (approximately) seventy internal commands of this form are no longer defined. If your class or package uses them then please replace them with new font commands described in L^{A}TEX 2ε Font Selection.

For example, the command \twlsf should be replaced by:

   \fontsize{12}{14}\normalfont\sffamily\selectfont

Another possibility is to use the rawfonts package, described in L^{A}TEX 2ε for Authors.

Also, remember that many of the fonts pre-loaded by L^{A}TEX 2.09 fonts are no longer pre-loaded.

\vpt \vipt \viipt . . .

These were the internal size-selecting commands in L^{A}TEX 2.09. (They can still be used in L^{A}TEX 2.09 compatibility mode.) Please use the command \fontsize instead: see L^{A}TEX 2ε Font Selection for details.

For example, \vpt should be replaced by:

   \fontsize{5}{6}\normalfont\selectfont
\prm, \pbf, \ppounds, \pLaTeX . . .

L^{A}TEX 2.09 used several commands beginning with \p in order to provide ‘protected’ commands. For example, \LaTeX was defined to be \protect\pLaTeX, and \pLaTeX was defined to produce the L^{A}TEX logo. This made \LaTeX robust, even though \pLaTeX was not.

These commands have now been reimplemented using \DeclareRobustCommand (described in Section 4.10). If your package redefined one of the \p-commands then you must remove the redefinition and use \DeclareRobustCommand to redefine the non-\p command.

\footheight \@maxsep \@dblmaxsep

These parameters are not used by L^{A}TEX 2ε so they have been removed, except in L^{A}TEX 2.09 compatibility mode. Classes should no longer set them.

References

[1] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The L^{A}TEX Companion. Addison-Wesley, Reading, Massachusetts, 1994.

[2] Donald E. Knuth. The TEXbook. Addison-Wesley, Reading, Massachusetts, 1986. Revised to cover TEX3, 1991.

[3] Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, Reading, Massachusetts, second edition, 1994.

L^{A}TEX 2ε Summary sheet: updating old styles

Section references below are to L^{A}TEX 2ε for Class and Package Writers.

1. Should it become a class or a package? See Section 2.3 for how to answer this question.

2. If it uses another style file, then you will need to obtain an updated version of this other file. Look at Section 2.7.1 for information on how to load other class and package files.

3. Try it: see Section 5.1.

4. It worked? Excellent, but there are probably still some things you should change in order to make your file into a well-structured L^{A}TEX 2ε file that is both robust and portable. So you should now read Section 2, especially 2.7. You will also find some useful examples in Section 3.

If your file sets up new fonts, font-changing commands or symbols, you should also read L^{A}TEX 2ε Font Selection.

5. It did not work? There are three possibilities here:

• error messages are produced whilst reading your file;

• error messages are produced whilst processing test documents;

• there are no errors but the output is not as it should be.

Don’t forget to check carefully for this last possibility.

If you have got to this stage then you will need to read Section 5 to find the solutions that will make your file work.